OpenStudy (anonymous):

csc^2 45degree- tan^2 45degree

OpenStudy (anonymous):

\[\csc ^{2}45degree= \left(\begin{matrix}\pi\\ 4\end{matrix}\right)=1\] \[\tan ^{2}=\left(\begin{matrix}\pi \\ 4\end{matrix}\right)=1\] I got zero and it doesnt look right for an answerr

OpenStudy (jdoe0001):

well, check your unit circle, what's the CSC(45)?

OpenStudy (anonymous):

\[\left(\begin{matrix} \pi \\4\end{matrix}\right)\]

OpenStudy (anonymous):

pi/4

OpenStudy (jdoe0001):

well, 45 degress expressed in radians is yes, pi/4, but that's not the value of it

OpenStudy (jdoe0001):

http://i.stack.imgur.com/r8uHr.gif :)

OpenStudy (anonymous):

\[\left[\begin{matrix}\sqrt{2} & 2 \\ \sqrt{2} & 2\end{matrix}\right]\]

OpenStudy (jdoe0001):

from the graph, those the (cos,sin) values, yes :)

OpenStudy (anonymous):

squr. root of 2 over 2 n sqru over 2. i subract both and i got zero

OpenStudy (jdoe0001):

well, you want ... CSC, no sine or cosine, check your trigonometry formulas cheatsheet for the "reciprocal identities", what is CSC equals to?

OpenStudy (anonymous):

\[\frac{ 1 }{ x}\]csc=

OpenStudy (jdoe0001):

http://www.sosmath.com/trig/Trig5/trig5/img1.gif

OpenStudy (anonymous):

i still lost

OpenStudy (anonymous):

1

OpenStudy (jdoe0001):

well, from the sheet on that graphic, you can see that \[csc(\theta)=\frac{1}{sin(\theta)}\] thus \[csc(45)=\frac{1}{sin(45)}\]

OpenStudy (jdoe0001):

\[sin(45)=\frac{\sqrt{2}}{2} \ and\ thus \frac{1}{\frac{\sqrt{2}}{2}}= \frac{2}{\sqrt{2}}=csc(45)\]